Eddie and Marco are working on a landing pad for their ultra light air craft that will be 10 feet longer than the width. If the area of the landing pad is to be 1200 square feet, find the length and the width.
(here's what I worked out and g got wrong: x=width in feet x+10 = length x^2 +10x+1200
x^2+10x+25=1200+25
(x+5)^2 +25 =1225
x+5 = plus or minus the square root of 1225 =35
x+5 = plus or minus 35
x= -5 plus or minus 35 ! ) help!

Answer by scott8148(6628)

you seem to have most of it ___ just a bit jumbled

area is 1200 ___ x^2 + 10x = 1200

completing the square ___ x^2 + 10x + 25 = 1225

taking square root ___ x + 5 = ± 35 ( you only need the ± on one side)

x + 5 = -35 ___ x = -40 ___ negative value not realistic

x + 5 = 35 ___ x = 30